OpenStudy (anonymous):

A 20 ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge.The horizontal distance that it spans is 4 ft longer than that the height reaches on the side of the bridge.Find the horizontal and vertical distances spanned by this brace.

OpenStudy (amistre64):

dude, I just figured out im am idiot :) 9 * 7 = 63 and 9 =2+7....

OpenStudy (anonymous):

what is this

OpenStudy (amistre64):

the last question you had; the area = 63; and length is w+2.....

OpenStudy (amistre64):

I missed it when I did (w+2)(w)... it equals w^2 + 2w.... NOT: w^2 + w my mistake

OpenStudy (anonymous):

you mean to tell me we did alllllllllll that for nothing...........it did seem a little complicated.....lol

OpenStudy (amistre64):

lol.....good practice I guess :)

OpenStudy (anonymous):

so our final answer is...........

OpenStudy (amistre64):

Lenght = 9 and width =7 is what we were supposed to get, and that i the correct answer.

OpenStudy (amistre64):

so this question here is a triangle: and we use the formula h^2 + b^2 = d^2 or whatever letter you wanna use. we know that the span is 4ft longer than its height

OpenStudy (amistre64):

lets set it up: h^2 + (h+4)^2 = 20^2 h^2 + h^2 +8h +16 = 400

OpenStudy (amistre64):

2h^2 +8h +16 -400 =0 2h^2 +8h -384 = 0 2(h^2 +4h -192) = 0...I hope im right so far :)

OpenStudy (amistre64):

2(0) = 0 so... h^2 +4h -192 = 0 h= -4/2 + sqrt(16 - (4)(1)(192))/2

OpenStudy (amistre64):

h= -2 + sqrt(784)/2

OpenStudy (amistre64):

h = 14 - 2 = 12 h=12 and span =16.

OpenStudy (amistre64):

when I double check, it works like a dream :)

OpenStudy (anonymous):

looks about right to me

OpenStudy (amistre64):

i had to get my brains back :)

OpenStudy (anonymous):

you been doing too much math.........lol

OpenStudy (amistre64):

lol.... i do agree

OpenStudy (anonymous):

I wanna be like you when I grow up..........lol!